YoursTruly:

Identify the slope and y-intercept of the equation -5x + 12y = 20. Show how you found them.

MARC:

Step 1:Arrange them into the slope-intercept form, y=mx+c

MARC:

y=mx+c where m=slope c=y-intercept

MARC:

add both sides by 5x

MARC:

\(\color{red}{5x}-5x+12y=\color{red}{5x}+20\)

Marziman:

Slope is 5/3 Intercept is -4

coacoapuffprincess:

Y cant u do these your self?

coacoapuffprincess:

I am

Korianna:

My bad I spelled ur name wrong

MARC:

@Marziman i did appreciate ur answer but it is not right to give a direct answer ...

coacoapuffprincess:

NO U DIDNT

MARC:

http://questioncove.com/rules.html @Marziman pls read this CoC Hv a nice day... Bye!

coacoapuffprincess:

thx

Marziman:

MARC Upon graphing the equation I was given that solution.